# Partially stochastic deep learning with uncertainty quantification for model predictive heating control

### v2

@@ 2.3 Deep learning models
This work focuses on supervised learning, in which models are trained to make predictions based on labeled input-output pairs: $\{\bm{X},\Delta\bm{y}\}$. Our inputs are defined as $\bm{X}\in\mathbb{R}^{N\times L\times M}$, where $N$ is the number of samples, $L$ is the sequence length (the number of preceding historical instances considered for the input variables given time $t$), and $M$ is the number of input variables. Similarly to our reference model, the input variables of the DL models implemented in this work (presented in Table 1) are designed to incorporate physics-based feature engineering. Therefore, instead of having one input dimension for each variable, some of the variables are combined to reflect their physical correlations related to the thermodynamical system: $T_{{sup}}-T_{{in}}$ and $\Phi_{{rad}}$ quantify the heat brought to the building by the heating system and solar irradiation, while $T_{{out}}-T_{{in}}$ describes the heat that leaks through the building’s envelope into the environment. The disturbance effect from internal heat sources is modeled by using a function $\xi$, which simply maps time $t$ to an integer value based on the hour of the week: 1–24 for non-business-day and 25–48 for business-day hours. Additionally, the azimuth and elevation angles of the sun, $\alpha_{azi}$ and $\alpha_{ele}$, are included, calculated based on the building’s GPS coordinates using the software package pvlib [34, 35]. These two additional variables enhance the modeling of the solar effect based on seasonality and building orientation. Finally, because DL models are capable of modeling complex non-linear relationships such as thermal lag effects, each input sequence matrix $\bm{X}^{(t)}=[\bm{x}^{(t-6)},\ldots,\bm{x}^{(t)}]\in\mathbb{R}^{L\times M}$ in $\bm{X}$ contains $L=7$ latest realizations of the $M=6$ input variables in Table 1. The length of the input sequence $L=7$ was selected based on experimenting with various values ranging from $1$ to $48$.
@@
Table 1: Input variables of the DL models.
To incorporate the physics of the modeled system given the input variables and their connection to (6), the DL models are trained to predict the change in indoor temperature $\Delta y^{(t)}=T_{{in}}^{(t+1)}-T_{{in}}^{(t)}$ instead of indoor temperature directly. Our previous experience indicates that with machine learning models, this approach results in more precise and physically meaningful predictions [9]. Thus, the outputs $\Delta\bm{y}\in\mathbb{R}^{N}$ represent the differences between the indoor temperature measurements of successive hours. As mentioned in 2.2, the time resolution of the observed data used in this work is hourly.
@@ 2.3.1 Deterministic deep learning model
Among NN architectures, RNNs are the most commonly used for modeling sequential data. They differ from standard feedforward networks in their ability to incorporate feedback loops, allowing them to utilize previous model states in their computations [36]. However, standard RNNs often face challenges such as vanishing or exploding gradients [37]. To address these issues, LSTMs were developed, incorporating internal gating mechanisms that allow the learning of long-term dependencies in sequential data [38]. LSTMs are specifically designed to mitigate the vanishing gradient problem by employing a technique called adaptive forgetting to reset model memory at appropriate time intervals [39].
@@
The structure of an LSTM layer consists of $L$ hidden LSTM units that are connected recursively. Each unit has a forget gate $\bm{f}^{(t)}$, an input gate $\bm{i}^{(t)}$, and a candidate memory gate $\bm{q}^{(t)}$, collectively responsible for managing the model memory, the cell state $\bm{c}^{(t)}$, by determining which stored information should be discarded and which new information should be inserted. The output gate $\bm{o}^{(t)}$ is needed when combining the information contained in the output (hidden state representation) of the previous LSTM unit $\bm{h}^{(t-1)}$, the current input to the unit $\bm{x}^{(t)}$, and the updated cell state $\bm{c}^{(t)}$. The mathematical formulas for the gate operations are given by [40]:
[…]
where $\sigma$ denotes the sigmoid activation function, $\bm{W}_{xi},\bm{W}_{xf},\bm{W}_{xo},\bm{W}_{hi},\bm{W}_{hf},\bm{W}_{ho}$ are weight parameters, $\bm{x}^{(t)}$ is the input data sequence, and $\bm{b}_{i},\bm{b}_{f},\bm{b}_{o}$ are bias parameters. Cell state $\bm{c}^{(t)}$ and hidden state $\bm{h}^{(t)}$ are then updated by:
$\begin{split}\bm{c}^{(t)}&=\bm{f}^{(t)}\bm{c}^{(t-1)}+\bm{i}^{(t)}\bm{q}^{(t)}\\ \bm{h}^{(t)}&=\bm{o}^{(t)}\tanh(\bm{c}^{(t)}),\end{split}$ (12)
where $\bm{W}_{xg}$ and $\bm{W}_{hg}$ are weight parameters, and $\bm{b}_{o}$ is a bias parameter.
We construct our deterministic DL model architecture as follows. An LSTM layer encodes the information in an input sequence $\bm{X}^{(t)}=[\bm{x}^{(t-(L-1))},\dots,\bm{x}^{(t)}]\in\mathbb{R}^{L\times M}$ into hidden states $\bm{h}^{(t-(L-1))},\dots,\bm{h}^{(t)}]\in\mathbb{R}^{L\times D_{h}}$. The combined information in the final hidden state representation $\bm{h}^{(t)}$ is decoded into a scalar prediction $\Delta\hat{y}^{(t)}$ by a multilayer perceptron (MLP) [36]. The MLP block is built from two linear layers connected by the ReLU activation function. The input dimensions of the linear layers are $D_{h}$ and $\frac{1}{2}D_{h}$, respectively. Based on previous experiments with different hidden dimensionalities, $D_{h}=2^{10}$ was selected [9]. This model architecture, which we refer to as LSTM+MLP, is visualized in Figure 2, and mathematically, we can write the forward action of the model given the inputs as
@@
where $\Psi^{\rm LSTM}(\bm{z}_{1})$ denotes the LSTM, and $\Psi^{\rm MLP}([\bm{z}_{2},\bm{z}_{3}])$ denotes the MLP computation. The parameters of the LSTM unit are denoted as $\bm{z}_{1}$, the parameters for the first linear layer are denoted as $\bm{z}_{2}$ and the parameters for the second linear layer used as the output layer are denoted as $\bm{z}_{3}$. Thus, all model parameters can be collectively expressed as $\bm{z}=[\bm{z}_{1},\bm{z}_{2},\bm{z}_{3}]$.
@@ 2.3.2 Partially stochastic deep learning model
Bayesian Neural Networks (BNNs) incorporate probabilistic modeling into conventional neural networks (NNs) by treating the parameters as random variables rather than fixed values. This allows for uncertainty quantification in predictions. The Bayesian aspect arises from assigning prior distributions to the weights and biases, and in this context, the training amounts to estimating the corresponding posterior distribution using Bayes’ theorem [41].
Not all parameters in a BNN need to be stochastic to achieve theoretical expressivity and gain in predictive performance [42]. With a smaller optimized parameter space, partially stochastic NNs can outperform fully stochastic NNs in training time. However, careful selection of the stochastic layer is crucial, as not all architectures may meet the criteria for effective probabilistic prediction. The idea is to split the parameter space into two sets, such that the parameter vector can be written as $\bm{z}=[\bm{z}_{D},\bm{Z}_{R}]$. Here, $\bm{z}_{D}$ represents the deterministic parameters for which point estimates are learned, while $\bm{Z}_{R}$ denotes the random parameters for which a distribution, or an approximation of it, is determined. The random parameters $\bm{Z}_{R}$ can be inferred jointly with the deterministic parameters or separately through a two-stage training process [42].
There are several alternatives on how to select the random parameters $\bm{Z}_{R}$. A common method involves isolating a specific layer and assigning all its parameters to $\bm{Z}_{R}$. Building upon the LSTM+MLP model (see 2.3.1) and experimenting with different layer parameterizations, we found that choosing the first linear layer as the stochastic layer yielded the best results in terms of predictive performance and computational cost. Thus, we propose a partially stochastic DL model architecture comprising an LSTM layer, a Bayesian linear layer, ReLU activation, and a standard linear layer. We refer to this model architecture as LSTM+BNN. Hence, the primary distinction between this model and the deterministic model described in Section 2.3.1 lies in the treatment of the first linear layer, which is now modeled as a Bayesian linear layer. In this case, the parameters $\bm{z}_{2}$ are represented as a random vector $\bm{Z}_{R}$. The parameters of the LSTM layer, $\bm{z}_{1}$, and the parameters in the second layer of the MLP, $\bm{z}_{3}$, are deterministic, so we lump them together as $\bm{z}_{D}=[\bm{z}_{1},\bm{z}_{3}]$. This generates a partially stochastic model of the form:
$\Psi(\bm{X};[\bm{Z}_{R},\bm{z}_{D}])=\left(\Psi^{\rm MLP}([\bm{Z}_{R},\bm{z}_{3}])\circ\Psi^{\rm LSTM}(\bm{z}_{1})\right)(\bm{X}).$ (15)
The training of the model (15) requires the estimation of the posterior distribution of the parameters $\bm{Z}_{R}$ given the data. This task is computationally intensive in high dimensions as we require MCMC methods to obtain samples from the posterior [41]. To maintain computational scalability, variational inference (see 2.1.1) is commonly performed, finding an approximating density to the posterior by maximizing the ELBO. Following (3), the ELBO for this task is expressed as:
@@
$\text{ELBO}(q)=\mathbb{E}_{q}\bigg{[}\left\|\Delta\bm{y}-\Psi(\bm{X};[\bm{Z}_{R},\bm{z}_{D}])]\right\|_{1}\bigg{]}-w_{\rm KL}D_{\rm KL}\left(q(\bm{Z}_{R})\parallel p(\bm{Z}_{R})\right),$ (16)
where $p(\bm{Z}_{R})$ is the prior distribution of the random parameters set as Gaussian $\mathcal{N}(\bm{0},\beta^{2}\bm{I})$. The likelihood term corresponds to a standard multivariate Laplace distribution, which is essentially the probabilistic version of the MAE loss in (14). A weighted KL divergence is employed to address the potential scale difference between the KL divergence term and the likelihood term, with a weight of $w_{\rm KL}=10^{-3}$ applied to effectively balance these terms. Note that the likelihood component ensures that the parameters fit the data effectively, while the KL divergence maintains consistency with the prior distribution.
The approximating density $q$ is assumed to be Gaussian $\mathcal{N}(\bm{\mu}_{R},\bm{\sigma}_{R}^{2}\bm{I})$, enabling the KL divergence term in (16) to be efficiently evaluated in closed form. During training, the optimization process iteratively updates $\bm{\mu}_{R},\bm{\sigma}_{R}$ (for the random parameters) and $\bm{z}_{D}$ (deterministic parameters) to minimize the negative ELBO. Specifically, at each forward pass through the model and for each sample in $\bm{X}$, the uncertain parameters $\bm{Z}_{R}$ are drawn from the current $\mathcal{N}(\bm{\mu}_{R},\bm{\sigma}_{R}^{2}\bm{I})$, and these values are incorporated into the likelihood component. Meanwhile, the deterministic parameters $\bm{z}_{D}$ are optimized directly during the training loop. This section concludes with a sketch of our approach in Algorithm 1, illustrating the main components of our model and its training process.
$\bm{X}\in\mathbb{R}^{N\times L\times M},\ \Delta\bm{y}\in\mathbb{R^{N}}$ $\triangleright$ Training data
$\bm{h}_{0}\in\mathbb{R}^{D_{h}}\leftarrow\bm{0}$ $\triangleright$ Initial LSTM hidden state
$\bm{c}_{0}\in\mathbb{R}^{D_{h}}\leftarrow\bm{0}$ $\triangleright$ Initial LSTM cell state
$p(\bm{Z}_{R})\leftarrow\mathcal{N}(\bm{0},\beta^{2}\bm{I})$ $\triangleright$ Prior of BNN parameters
$w_{\rm KL}\leftarrow 10^{-3}$ $\triangleright$ Define weight for KL divergence
Initialize model parameters $\bm{\mu}_{R},\bm{\sigma}_{R},\bm{z}_{D}$
1Initialize Minimum validation loss $\leftarrow\infty$
2procedure Training($\bm{X}$)
3  $\Delta\bm{y}=\{\Delta\bm{y}_{\text{train}}\in\mathbb{R}^{N_{\text{train}}},\Delta\bm{y}_{\text{val}}\in\mathbb{R}^{N_{\text{val}}}\}$ for $i=1$ to epochs do
4
5 end for
Draw $\bm{Z}_{R}\sim q(\bm{Z}_{R})$
[…]
  Back propagate loss to compute gradients w.r.t. $\bm{\mu}_{R},\bm{\sigma}_{R},\bm{z}_{D}$
  Update $\bm{\mu}_{R},\bm{\sigma}_{R},\bm{z}_{D}$ based on learning rate and gradients $\triangleright$ Adam optimizer
  $q(\bm{Z}_{R})\leftarrow\mathcal{N}(\bm{\mu}_{R},\bm{\sigma}^{2}_{R}\bm{I})$
  Draw $\bm{Z}_{R}\sim q(\bm{Z}_{R})$
[…]
7if  Validation loss << Minimum validation loss then
8
9 end if
Save LSTM+BNN model with current parameters $\bm{\mu}_{R},\bm{\sigma}_{R},\bm{z}_{D}$ as the Best model
10  Minimum validation loss $\leftarrow$ Validation loss
11  return Best model
Algorithm 1 LSTM+BNN: forward action and training sketch.
@@ 3.1 Dataset
The building-specific datasets were collected by a private company specializing in MPC for centrally-heated residential buildings. For a total of $S=100$ buildings, the datasets contain hourly time series data between years 2017 and 2022. The indoor temperature data represent the averages of sensor measurements from each apartment in the buildings. The space heating supply water temperature data are measured at the building-specific heating substations. The outdoor temperature and the solar irradiation data are from the professional weather data provider of the company, fetched for the buildings based on their GPS coordinates.
Since the heating season in northern climates spans September to May, only data from these months are considered. During the summer months, the heating systems are shut off, and therefore the indoor temperature cannot be affected by controlling the supply water temperature of the space heating circuit.

### v3

@@ 2.3 Deep learning models
This work focuses on supervised learning, in which models are trained to make predictions based on labeled input-output pairs: $\{\bm{X},\Delta\bm{y}\}$. Our inputs are defined as $\bm{X}\in\mathbb{R}^{N\times L\times M}$, where $N$ is the number of samples, $L$ is the sequence length (the number of preceding historical instances considered for the input variables given time $t$), and $M$ is the number of input variables. Similarly to our reference model, the input variables of the DL models implemented in this work (presented in Table 1) are designed to incorporate physics-based feature engineering. Therefore, instead of dedicating one input dimension to each raw variable, some of the variables are combined to reflect their physical correlations related to the thermodynamic system: $T_{{sup}}-T_{{in}}$ and $\Phi_{{rad}}$ quantify the heat brought to the building by the heating system and solar irradiation, while $T_{{out}}-T_{{in}}$ describes the heat that leaks through the building’s envelope into the environment. The disturbance effect from internal heat sources is modeled by using a function $\xi$, which simply maps time $t$ to an integer value based on the hour of the week: 1–24 for non-business-day and 25–48 for business-day hours. Additionally, the azimuth and elevation angles of the sun, $\alpha_{azi}$ and $\alpha_{ele}$, are included, calculated based on the building’s GPS coordinates using the software package pvlib [45, 46]. These two additional variables enhance the modeling of the solar effect based on seasonality and building orientation. Finally, to leverage the ability of DL models to capture complex non-linear relationships such as thermal lag effects, each input sequence matrix $\bm{X}^{(t)}=[\bm{x}^{(t-6)},\ldots,\bm{x}^{(t)}]\in\mathbb{R}^{L\times M}$ in $\bm{X}$ contains $L=7$ most recent hourly observations of the $M=6$ input variables listed in Table 1. The length of the input sequence $L=7$ was selected based on experiments with various values ranging from $1$ to $48$.
Table 1: Input variables of the DL models.
@@
To incorporate the physics of the modeled system given the input variables and their connection to (6), the DL models are trained to predict the change in indoor temperature $\Delta y^{(t)}=T_{{in}}^{(t+1)}-T_{{in}}^{(t)}$ instead of indoor temperature directly. Our previous experience indicates that for machine learning models, this approach yields more precise and physically meaningful predictions [9]. Thus, the outputs $\Delta\bm{y}\in\mathbb{R}^{N}$ represent the differences between the indoor temperature measurements of successive hours. As mentioned in Section 2.2, the time resolution of the observed data used in this work is hourly.
@@ 2.3.1 Deterministic deep learning model
Among neural network architectures, RNNs are the most commonly used for modeling sequential data. They differ from standard feedforward networks in their ability to incorporate feedback loops, allowing them to utilize previous model states in their computations [47]. However, standard RNNs often face challenges such as vanishing or exploding gradients [48]. To address these issues, LSTMs were developed, incorporating internal gating mechanisms that allow the learning of long-term dependencies in sequential data [49]. LSTMs are specifically designed to mitigate the vanishing gradient problem by employing a technique called adaptive forgetting to reset model memory at appropriate time intervals [50].
The structure of an LSTM layer consists of $L$ hidden LSTM units that are connected recursively. Each unit has a forget gate $\bm{f}^{(t)}$, an input gate $\bm{i}^{(t)}$, and a candidate memory gate $\bm{q}^{(t)}$, collectively responsible for managing the model memory, the cell state $\bm{c}^{(t)}$, by determining which stored information should be discarded and which new information should be inserted. The output gate $\bm{o}^{(t)}$ is needed when combining the information contained in the output (hidden state representation) of the previous LSTM unit $\bm{h}^{(t-1)}$, the current input to the unit $\bm{x}^{(t)}$, and the updated cell state $\bm{c}^{(t)}$. The mathematical formulas for the gate operations are given as:
[…]
where $\sigma$ denotes the sigmoid activation function, $\bm{W}_{xf},\bm{W}_{xi},\bm{W}_{xq},\bm{W}_{xo},\bm{W}_{hf},\bm{W}_{hi},\bm{W}_{hq},\bm{W}_{ho}$ are weight parameters, and $\bm{b}_{f},\bm{b}_{i},\bm{b}_{o},\bm{b}_{q}$ are bias parameters. Cell state $\bm{c}^{(t)}$ and hidden state $\bm{h}^{(t)}$ are then updated by:
$\begin{split}\bm{c}^{(t)}&=\bm{f}^{(t)}\bm{c}^{(t-1)}+\bm{i}^{(t)}\bm{q}^{(t)}\\ \bm{h}^{(t)}&=\bm{o}^{(t)}\tanh(\bm{c}^{(t)}).\end{split}$ (12)
We construct our deterministic DL model architecture as follows. An LSTM layer encodes the information in an input sequence $\bm{X}^{(t)}=[\bm{x}^{(t-(L-1))},\dots,\bm{x}^{(t)}]\in\mathbb{R}^{L\times M}$ into hidden states $\bm{h}^{(t-(L-1))},\dots,\bm{h}^{(t)}]\in\mathbb{R}^{L\times D_{h}}$. The combined information in the final hidden state representation $\bm{h}^{(t)}$ is decoded into a scalar prediction $\Delta\hat{y}^{(t)}$ by a multilayer perceptron (MLP) [47]. The MLP block is built from two linear layers connected by the ReLU activation function. The first linear layer has an input dimension of $D_{h}$, and the second (output) layer has an input dimension of $\frac{1}{2}D_{h}$. Based on previous experiments with different hidden dimensionalities, $D_{h}=2^{10}$ was selected [9]. This model architecture, which we refer to as LSTM+MLP, is visualized in Figure 2.
@@
Mathematically, we can write the forward action of the model given the inputs as
@@
where $\Psi^{\rm LSTM}(\bm{z}_{1})$ denotes the LSTM, and $\Psi^{\rm MLP}(\bm{z}_{2},\bm{z}_{3})$ denotes the MLP computation. The parameters of the LSTM layer are denoted as $\bm{z}_{1}$, the parameters for the first linear layer are denoted as $\bm{z}_{2}$, and the parameters for the second linear layer used as the output layer are denoted as $\bm{z}_{3}$. Thus, all model parameters can be collectively expressed as $\bm{z}=[\bm{z}_{1},\bm{z}_{2},\bm{z}_{3}]$.
@@ 2.3.2 Partially stochastic deep learning model
Bayesian Neural Networks (BNNs) extend conventional neural networks by treating the model parameters (weights and biases) as random variables rather than fixed values. By placing prior distributions on these parameters and using Bayes’ theorem (1) to compute posterior distributions given data, BNNs naturally provide UQ for both their parameters and predictions [51].
Making all parameters in a DL model stochastic significantly increases the computational cost due to the higher dimensionality of the optimized parameter space. However, research by Sharma et al. [52] shows that expressive predictive distributions can be achieved through partially stochastic networks, where only a subset of parameters are treated probabilistically. This approach maintains the benefits of UQ while being more computationally efficient.
Nevertheless, careful selection of the stochastic parameters is crucial, as not all architectures may meet the criteria for effective probabilistic prediction. There are several alternatives on how to do this selection, and a common method involves isolating a specific layer and treating all its parameters as stochastic.
Our partially stochastic architecture, which we call LSTM+BNN, builds upon the deterministic LSTM+MLP model (see Section 2.3.1). Through experimentation with different layer configurations, we found that making only the first linear layer stochastic provided the best balance of predictive performance, expressivity, and computational efficiency. Thus, in our LSTM+BNN architecture, the first linear layer parameters ($\bm{z}_{2}$) are treated as stochastic, denoted as $\bm{Z}_{R}$. The LSTM and the second linear layer parameters remain deterministic, collectively denoted as $\bm{z}_{D}=[\bm{z}_{1},\bm{z}_{3}]$. The resulting partially stochastic model has the form:
$\Psi(\bm{X};[\bm{Z}_{R},\bm{z}_{D}])=\left(\Psi^{\rm LSTM}(\bm{z}_{1})\circ\Psi^{\rm MLP}([\bm{Z}_{R},\bm{z}_{3}])\right)(\bm{X}).$ (15)
The training of the model (15) requires the estimation of the posterior distribution for the random parameters $\bm{Z}_{R}$. However, computing the exact posterior in this case is intractable. Following the variational inference approach (described in Section 2.1.1), we approximate the true posterior with a tractable density $q(\bm{z}_{R})=\mathcal{N}(\bm{\mu}_{R},\bm{\sigma}_{R}^{2}\bm{I})$, parameterized by mean $\bm{\mu}_{R}$ and variance $\bm{\sigma}_{R}^{2}$. We maximize the ELBO, which can be written for our model as:
@@
$\text{ELBO}(q)=\mathbb{E}_{q}\bigg[\left\|\Delta\bm{y}-\Psi(\bm{X};[\bm{Z}_{R},\bm{z}_{D}])\right\|_{1}\bigg]-w_{\rm KL}D_{\rm KL}\left(q(\bm{z}_{R})\parallel p(\bm{z}_{R})\right),$ (16)
where $p(\bm{z}_{R})=\mathcal{N}(\bm{0},\beta^{2}\bm{I})$ is the Gaussian prior density for the random parameters $\bm{Z}_{R}$. The first term is the expected loss under $q(\bm{z}_{R})$, using the L1-norm (equivalent to a multivariate Laplace likelihood) to maintain consistency with the MAE loss (14) used in LSTM+MLP. The second term is the weighted negative KL divergence between the approximate density $q(\bm{z}_{R})$ and the prior $p(\bm{z}_{R})$. Since both densities are Gaussian, they form a conjugate exponential family pair, enabling the KL divergence to be evaluated efficiently in closed form. The weight $w_{\rm KL}=10^{-3}$ balances the scale difference between these two terms in (16). Intuitively, the first term ensures that the model fits the data effectively, while the second term maintains consistency with our prior beliefs.
During training, the variational parameters $\bm{\mu}_{R}$, $\bm{\sigma}_{R}$, and deterministic parameters $\bm{z}_{D}$ are optimized simultaneously. For a detailed description of the LSTM+BNN training procedure, see Algorithm 1 in A.
To generate probabilistic predictions from the LSTM+BNN model, the following Monte Carlo sampling strategy is applied. An input sequence $\bm{X}^{(t)}$ is passed through the model in (15) multiple times. At each forward pass, a realization $\bm{z}_{R}$ of the random parameters $\bm{Z}_{R}$ is sampled from the approximate posterior distribution $q(\bm{z}_{R})$ parameterized by the optimized $\bm{\mu}_{R}$ and $\bm{\sigma}_{R}$. This yields an array of predictions, from which the predictive mean and the standard deviation—a measure of predictive uncertainty—can be calculated.
@@ 3.1 Dataset
The building-specific datasets were collected by a private company specializing in MPC for centrally-heated residential buildings. For a total of $S=100$ buildings, the datasets contain hourly time series data between the years 2017 and 2022. The indoor temperature data represent the averages of sensor measurements from apartments within each building. The space heating supply water temperature data are measured at the building-specific heating substations. The outdoor temperature and the solar irradiation data are from the professional weather data provider of the company, fetched for the buildings based on their GPS coordinates. The availability of all these data is an essential prerequisite for the practical implementation of the indoor temperature models covered in this work.
Since the heating season in northern climates spans September to May, only data from these months are considered. During the summer months, the heating systems are shut off, and therefore, the indoor temperature cannot be affected by controlling the supply water temperature of the space heating circuit.